9 A manufacturer claims that 10% of women using the "pill" suffer from side effects. The...

9

A manufacturer claims that 10% of women using the "pill" suffer from side effects. The Federal Drug Administration (FDA) believes that the manufacturer's claim is too low and decides to test the manufacturer's claim at α = 5 %. A random sample of 900 women who use the manufacturer's pill shows that 103 suffer side effects.

The sample proportion of women who suffer effects from the pill is:
Answer #1

9)

Solution :

The sample proportion of women who suffer effects from the pill is: x / n = 103 / 900 = 0.1144